PSfrag replacements 2 3

243A: y^{2} + y = x^{3} − 1 243B: y^{2} + y = x^{3} + 2

243C: surface 243D: surface 243E: three-fold 243F: three-fold

images of 81A: four-fold

images of 27A (y^{2} + y = x^{3} − 7): three-fold

The Jacobian of X_{0}(243)